A material ‘B’ has twice the specific resistance of ‘A’. A circular wire made of ‘B’ has twice the diameter of a wire made of ‘A’. Then for the two wires to have the same resistance, the ratio AA /AB of their respective lengths must be ?

Grade:12

1 Answers

SHAIK AASIF AHAMED
9 years ago
Hello student,
R1=\frac{\rho _{a}l_{a}}{\pi R_{a}^{2}}
R2=\frac{\rho _{b}l_{b}}{\pi R_{b}^{2}}
so we get\frac{l_{b}}{l_{a}}=2
Thanks and Regards
Shaik Aasif
askIITians faculty